Solve xp2-yp-y=0


1
Expert's answer
2021-06-20T18:49:46-0400

Solve


"xp^2-yp-y=0, p=\dfrac{dy}{dx}=y'"

Solve for "y"


"y=\dfrac{xp^2}{p+1}"

Differentiate both sides with respect to "x"


"y'=\dfrac{(p^2+2xpp')(p+1)-xp^2p'}{(p+1)^2}"

"p^3+2p^2+p=p^3+p^2+xpp'(2p+2-p)"

"p(p+1)=xpp'(p+2)"

If "p=0"


"x(0)^2-y(0)-y=0=>y=0"

If "p\not=0"


"p+1=p'(p+2)x"


"p'(\dfrac{p+2}{p+1})=\dfrac{1}{x}"

"dp(1+\dfrac{1}{p+1})=dx(\dfrac{1}{x})"

"p+\ln|p+1|=\ln|x|-\ln C"

"x=C(p+1)e^p"

"p=\dfrac{y\pm\sqrt{y^2+4xy}}{2x}"


"x=C(\dfrac{y\pm\sqrt{y^2+4xy}}{2x}+1)e^{({y\pm\sqrt{y^2+4xy} \over 2x})}"